# Path Planning for a Cooperative Navigation Aid Vehicle to Assist Multiple Agents Intermittently

### v1

@@ 4 Intermittent Aiding Algorithm
        $\Delta^{\rm max}_{i}$ $=[J_{i}^{\rm max}-J_{i}(Z_{\rm cand})]/J_{i}^{\rm max}$ (44)
        $\Delta^{\rm opt}_{i}$ $=[J_{i}^{*}(k)-J_{i}(Z_{\rm cand})]/J_{i}(Z_{i}^{\rm cand})\;.$ (45)
@@
    If $D=N+1$ then the current task sequence ${A}$ has $N$ elements and is missing one task. Compare inserting the surfacing task at each position in ${A}$ and return the task sequence ${A}^{*}$ that yields the lowest cost (32).
@@
    If $D\leq N$ then the current task sequence ${A}$ already has $D$ agent tasks. Compare replacing each agent task with a surfacing task and return the task sequence ${A}^{*}$ that yields the lowest cost (32).
The algorithm described above is a greedy approach that selects the highest reward at each step in the task allocation. The reward (43) was designed to balance three objectives: (1) to secure a large reduction in the overall cost by selecting a large $\Delta^{\rm max}_{i}$, (2) to avoid a penalty for aiding an agent at a sub-optimal time (i.e., far away from its optimal time-to-aid as quantified by $\Delta^{\rm opt}_{i}$), and (3) to avoid long transit times $\tau_{i}$. Each of the three terms in (43) have similar magnitude and are scaled by user-selected weights $(\alpha,\beta,\gamma)$.
The greedy heuristic algorithm described above is efficient to compute, but it is suboptimal. For a sufficiently $N$, all of the candidate task sequences in $\mathcal{F}$ can be enumerated and exhaustively compared to find the optimal task sequence — we refer to this approach as the optimal (exhaustive enumeration) algorithm.
@@ 5.1 Illustrative Example
Table 1: Table of parameters used in numerical simulations. LU = length unit, and TU = time unit.
@@
[…]
Numerical simulations display the 50% confidence ellipse for the uncertain position of the agents. The CNA path was solved for using the optimal (exhaustive enumeration) algorithm to give the task sequence ${A}=\{2,0,1,3,4\}$. That is, the CNA first aids agent 2 (orange), then surfaces, then continues to aid agents 1 (green), 3 (blue), and 4 (purple). The overall cost of this task sequence was $J=1172$. Even though the agent 1 had the largest initial variance, the optimal task is to aid agent 2 first. Intuitively, agent 2 is closer to the CNA starting location and it more efficient to aid them along the way to agent 1. The CNA reached the final agent 4 at time $k=1233$. A modified problem formulation that allows the CNA to revisit agents with remaining mission time may be considered in future work.
@@ 5.2 Monte Carlo Simulation
The greedy algorithm was tested through randomized trials that varied the initial locations and heading angles of the agents, along with their initial variance. The parameters in the simulation are listed in Table 1. The agent initial positions and headings were randomized according to the following three strategies:
@@
The number of agents was varied from $N=3$ to $N=14$ and for each $N$ value 100 random scenarios were generated. Approximately 1/3 of each set of trials adopted each of the three initialization strategies described above and the initial variance of the agents was randomly chosen from the interval $\nu_{0|0}^{i}\in[0,3000]$. The greedy heuristic search algorithms were evaluated with the following combination of weights: G1 $=(\alpha,\beta,\gamma)=(1,0,0)$ , G2 $=(\alpha,\beta,\gamma)=(0,1,0)$ , G3 $=(\alpha,\beta,\gamma)=(0,0,1)$ , and G4 $=(\alpha,\beta,\gamma)=(1,0.5,0.5)$. This choice of algorithm parameters allows comparing the influence of each term in the reward (43), as well as their combined effect. With $N=10$ agents each trial took about 10 minutes to solve using the optimal algorithm, thus evaluations with this algorithm were limited to $N\leq 10$. Since the optimal algorithm tests all possible task sequences, the worst-case choice of task sequence was also recorded for comparison. The algorithms were all implemented in MATLAB and evaluated using a parallelized computation on a laptop computer.
The mean cost and the mean computation time for each algorithm as a function of number of agents is shown in Fig. 4. By comparing the optimal and worst-case performance to the bounds of (41), it is apparent that the bounds bracket the performance but are not tight. When comparing the heuristic algorithms, the G4 set of $(\alpha,\beta,\gamma)$ parameters that mix all three terms in (43) outperforms G1, G2, G3, which each only use one of the terms in (43). While the computation time of the optimal (exhaustive) algorithm increases exponentially, the average computation time of the greedy algorithms remains below $0.01$ seconds. Interestingly, the computation time exhibits an intuitive increase for a smaller agents — this may be due to how the parallel computation was handled by MATLAB.
@@ 6 Conclusion
This paper considered the problem of planning a planar path for a CNA to intermittently provide navigation aiding measurements to a set of $N$ agents while also considering the possibility of surfacing to improve its own navigation uncertainty. A greedy algorithm was formulated to solve for the desired CNA path by considering a heuristic that considers three weighted factors: (1) selecting agents that lead to large reductions in cost, (2) selecting agents for whom the aiding time is close to the optimal time-to-aid, and (3) selecting agents that are close to the CNA. A Monte Carlo experiment comparing the greedy algorithm to an optimal (exhaustive enumeration) algorithm showed that by combining all three weighted factors a lower cost path is returned than by considering each factor independently.
Future work may consider a CNA that can revisit agents in the task sequence or surface multiple times. The problem can also be extended to incorporate more sophisticated vehicle motion and sensor models and a finite communication radius that allow multiple measurements to be exchanged per intertask. Aiding agents that drift significantly from their nominal trajectories is also of interest.

### v2

@@ 4 Sequential Aiding Algorithm
        $\Delta^{\rm max}_{i}$ $=[J_{i}^{\rm max}-J_{i}(Z_{\rm cand})]/J_{i}^{\rm max}$
        $\Delta^{\rm opt}_{i}$ $=[J_{i}^{*}(k)-J_{i}(Z_{\rm cand})]/J_{i}(Z_{i}^{\rm cand})\;.$
@@
    If $D=N+1$ then the current task sequence ${A}$ has $N$ elements and is missing one task. Compare inserting the surfacing task at each position in ${A}$ and return the lowest cost sequence ${A}^{*}$ according to (23).
@@
    If $D\leq N$ then the current task sequence ${A}$ cannot be enlarged. Compare replacing each agent task with a surfacing task and return the task sequence ${A}^{*}$ that yields the lowest cost (23).
The above greedy algorithm selects the highest reward at each step in the task allocation. The reward (29) is designed to balance three objectives: (1) to secure a large reduction in the overall cost by selecting a large $\Delta^{\rm max}_{i}$, (2) to avoid a penalty for aiding an agent at a sub-optimal time (i.e., far away from its optimal time-to-aid as quantified by $\Delta^{\rm opt}_{i}$), and (3) to avoid long transit times $\tau_{i}$. Each of the three terms in (29) have similar magnitude and are scaled by user-selected weights $(\alpha,\beta,\gamma)$.
The greedy heuristic algorithm described above is efficient to compute, but is suboptimal. For a sufficiently large $N$, all of the candidate task sequences in $\mathcal{F}$ can be enumerated and exhaustively compared to find the optimal task sequence — we refer to this approach as the optimal (exhaustive enumeration) algorithm.
@@ 5.1 Illustrative Example
Table 1: Parameters used in numerical simulations. LU = length unit, TU = time unit.
@@
Process variance of agents $\nu_{w}$ 1 LU2
Process variance of CNA $\nu_{c}$ $\nu_{w}/10$ LU2
Initial variance of agents $\nu^{i}_{0|0}$ LU2
Measurement noise $\nu_{y}$ 10 LU2
Variance of CNA upon GPS update $\nu_{G}$ 10 LU2
Numerical simulations display the 50% confidence ellipse for the uncertain position of the agents. The CNA path was solved for using the optimal (exhaustive enumeration) algorithm to give the task sequence ${A}=\{2,0,1,3,4\}$. That is, the CNA first aids agent 2 (orange), then surfaces, then continues to aid agents 1 (green), 3 (blue), and 4 (purple). The overall cost of this task sequence was $J=1172$. Even though the agent 1 had the largest initial variance, the algorithm chooses to aid agent 2 first. Intuitively, agent 2 is closer to the CNA starting location and it more efficient to aid them along the way to agent 1. The CNA reached the final agent 4 at time $k=1233$. A modified problem formulation that allows the CNA to revisit agents with remaining mission time may be considered in future work.
@@ 5.2 Monte Carlo Simulation
The greedy algorithm was tested through a Monte Carlo simulation that randomized the initial locations, heading angles, and initial variance of the agents according to the following three strategies:
@@
The number of agents was varied from $N=3$ to $N=14$ and for each $N$ value 100 random scenarios were generated. The parameters in the simulation are listed in Table 1. Approximately 1/3 of each set of trials adopted each of the three initialization strategies described above and the initial variance of the agents was randomly chosen from the interval $\nu_{0|0}^{i}\in(0,3000]$. The greedy heuristic search algorithms were evaluated with the following combination of weights: G1 $=(\alpha,\beta,\gamma)=(1,0,0)$ , G2 $=(\alpha,\beta,\gamma)=(0,1,0)$ , G3 $=(\alpha,\beta,\gamma)=(0,0,1)$ , and G4 $=(\alpha,\beta,\gamma)=(1,0.5,0.5)$. This allowed comparing the influence of each term in the reward (29), as well as their combined effect. With $N=10$ agents each trial took about 10 minutes to solve using the optimal algorithm, thus evaluations with this algorithm were limited to $N\leq 10$. Since the optimal algorithm tests all possible task sequences, the worst-case choice of task sequence was also recorded for comparison. The algorithms were all implemented in MATLAB using a parallelized computation on a laptop computer.
The mean cost and the mean computation time for each algorithm is shown in Fig. 4. By comparing the optimal and worst-case performance to the bounds of (27), it is apparent that the bounds bracket the performance but are not tight. When comparing the heuristic algorithms, the G4 set of $(\alpha,\beta,\gamma)$ parameters that mix all three terms in (29) outperforms G1, G2, G3, which each only use one of the terms in (29). While the computation time of the optimal (exhaustive) algorithm increases exponentially, the average computation time of the greedy algorithms remains below $0.01$ seconds. Unexpectedly, the computation time increases slightly for a smaller number of agents — this may be due to how the experiment was implemented with the parallel processing toolbox in MATLAB.
@@ 6 Conclusion
This paper considered the problem of planning a planar path for a CNA to sequentially provide navigation aiding measurements to a set of $N$ agents while also considering the possibility of surfacing to improve its own navigation uncertainty. A greedy algorithm was formulated to solve for the desired CNA path by considering a heuristic that considers three weighted factors: (1) selecting agents that lead to large reductions in cost, (2) selecting agents for whom the aiding time is close to the optimal time-to-aid, and (3) selecting agents that are close to the CNA. A Monte Carlo experiment comparing the greedy algorithm to an optimal algorithm showed that by combining all three weighted factors a lower cost path is returned than by considering each factor independently.
Future work may consider heuristics with recipes for choosing various tuning parameters or other optimization schemes. Other scenarios could also be investigated—for example, a CNA that can revisit agents and surface multiple times, a CNA that provides water current estimates to agents to minimize final position errors, or a CNA with a finite communication radius that can service multiple agents simulatenously. Aiding agents with unknown trajectories (i.e., requiring an initial search to localize them) may also be of interest.